Safawi kicked a ball with an initial speed of 20 ms-1 at 30º angle. If the ball experienced a constant vertical acceleration of -9.81 ms-2, calculate the ball’s maximum height and distance. Explain how soccer player could apply the force of summation theory to optimize his kicks performance. 


1
Expert's answer
2022-02-08T10:12:26-0500

Given:

"v_0=20\:\rm m/s"

"\theta=30^\circ"

"g=9.81\:\rm m/s^2"


The maximum height

"h_{\max}=\frac{v_0^2\sin^2\theta}{2g}=\frac{20^2\sin^230^\circ}{2*9.81}=5.1\:\rm m"

The range


"L_{\max}=\frac{v_0^2\sin2\theta}{g}=\frac{20^2\sin60^\circ}{9.81}=35.3\:\rm m"